OpenStudy (anonymous):

Find the solution to the equation 64^(3 – x) = 4^(2x). is the answer x=9/5 ?

OpenStudy (anonymous):

\[64^{\left( 3-x \right)}=\left( 4^3 \right)^{\left( 3-x \right)}=4^{3\left( 3-x \right)}=4^{9-3x}\] now you can solve.

OpenStudy (anonymous):

i think you are correct.

OpenStudy (anonymous):

oh ok:)

OpenStudy (anonymous):

yw